OpenStudy (anonymous):

solve by the elimination method 7r-8s=55 8r+7s=79

hero (hero):

What the...?

hero (hero):

I know I helped you with this before

hero (hero):

What happened?

OpenStudy (anonymous):

I only solved for r

hero (hero):

to solve for s, replace what you found for r into the equation

OpenStudy (anonymous):

multiply first line by 8 and second line by -7. Sum both. 56r -64s=440 -56r -49s=-553 ---------------- 0 -113s=-113 s=1 subst on first line: r=9

OpenStudy (anonymous):

Thanks Leo , that was crystal clear.

hero (hero):

Yeah, but...

OpenStudy (anonymous):

the squeaky wheel gets the greese, and other army maxims....

hero (hero):

Leo's method doesn't allow you to play with the numbers as much

OpenStudy (anonymous):

Hero, i solve by elimination method... i eliminate the r and found s.

hero (hero):

I eliminated s and found r.

hero (hero):

You do realize that r comes before s. You're supposed to find that first.

hero (hero):

Think x and y

OpenStudy (anonymous):

its the same!

hero (hero):

Not really. A true mathematician follow the correct order and procedure of things.

hero (hero):

If I was your teacher I would take a quarter of a point off for each wrong step

OpenStudy (anonymous):

ok. Im not 'true mathematician'. Im engineer.. i solve problems. =)

OpenStudy (anonymous):

You right.

hero (hero):

If you're an engineer, I hope you're following the instructions in the right order.